# Rajatsubhra Chakraborty Arkaprava Sinha  Dominick Reilly  Manish Kumar Govind Pu WangFrancois BremondSrijan Das UNC Charlotte  Inria  Université Côte d’Azur

### v1

@@ 3.3 Action-Conditioned Object Cue for LLAVIDAL
Action conditioned object detection. Given a stitched ADL video, which comprises a sequence of trimmed video segments (denoted as $clip_{j}$), the first stage extracts the categories of objects present that are pertinent to the actions performed within each clip. We uniformly sample 8 frames from each video and employ a pre-trained BLIP-2 model [45] to generate a list of distinct objects observed in the frames. To avoid training LLAVIDAL with noisy data, we perform a filtering on the list of objects using the ground-truth action labels and GPT-3.5. Specifically, for each $clip_{j}$ within a stitched video, we input the corresponding action label and the list of detected objects to GPT-3.5 and prompt it to identify the object(s) most relevant to the given action. For instance, if the objects plant, chair, bottle, table are detected in a video labeled with the action Drinking, GPT-3.5 is expected to filter out and select [bottle] as the relevant object for $clip_{j}$. Refer to Appendix H for our detailed action conditioned object detection prompting strategy.
Object Localization and Tracking. Given the list of relevant objects identified in the first stage, the second stage involves spatial localization of these objects within the scene and their temporal association (i.e., object tracking) based on the feature similarity of the image regions corresponding to the localized objects in the stitched video. We employ a pre-trained open vocabulary object localization model (ObjectLM), OWLv2 [50], and input the list of relevant objects detected in stage 1 along with the corresponding video. Localization and tracking are performed on 8 frames that are uniformly sampled from $clip_{j}$ within a stitched video. For each frame, we obtain bounding boxes $B_{t}\in\mathbb{R}^{n\times 4}$, where each bounding box corresponds to one of the $n$ relevant objects in the $t$th frame. Features for each object are then extracted from the image regions within these bounding boxes using our object localization model. We denote the features for the objects in frame $t$ as $O_{t}\in\mathbb{R}^{8n\times D_{o}}$, where $D_{o}$ is the object feature dimension. To associate objects across frames, we utilize a feature-based object tracking approach. Specifically, for each object in frame $t$, represented by the feature vector $O_{i}^{t}\in\mathbb{R}^{D_{o}}$, we compute the cosine similarity between $O_{i}^{t}$ and all feature vectors in frame $t+1$. The object $i$ in frame $t$ is then associated with the object in frame $t+1$ that exhibits the highest similarity score. This matching process is iterated for all objects in each frame, thereby establishing a track for each relevant object throughout the sampled frames. These object tracks, with corresponding bounding boxes and features, facilitate the integration of object information into the training of LLAVIDAL: Object as QA, Object as context, and Object as features.
Object as QA. Similar to the approach taken with poses, to generate QA pairs for objects, we formulate a question based on the trajectory coordinates of the relevant object(s). These QA pairs are added to the set of text queries $Q_{t}$ for instruction tuning LLAVIDAL.
Object as Context. To integrate the context of detected objects into the LLM space, we append the list of relevant object labels, denoted by $Q^{o}_{t}$, to each text query token $Q_{t}$. Consequently, the updated text query is represented as $Q_{t}^{new}=[Q^{o}_{t}\hskip 3.61371ptQ_{t}]$. This enhanced text query, $Q_{t}^{new}$, is utilized for instruction tuning.
Object as Features. The object features extracted during the object localization and tracking stage are utilized as input tokens $Q_{o}\in\mathbb{R}^{8n\times D_{o}}$, which are incorporated alongside the text query tokens ($Q_{t}$) and input video tokens ($Q_{v}$). For $n$ relevant objects detected, the object query $Q_{o}$ is structured using the following template $[\langle Q_{o}\rangle\hskip 3.61371pt=\langle Q_{o}^{1}\rangle\hskip 3.61371pt\langle Q_{o}^{2}\rangle\hskip 3.61371pt...\langle Q_{o}^{n}\rangle]$ where $Q_{o}^{j}\in\mathbb{R}^{8\times D_{o}}$ represent the features of each relevant object in the video.
3.4 Training LLAVIDAL
As illustrated in Figure 3, the QA pairs, along with context or features obtained from the RGB video, 3D poses, and object cues can be integrated into LLAVIDAL. Integrating QA pairs and contextual information is straightforward; they are introduced into $Q_{t}$ and trained using standard methods for LLVM. However, to integrate other modalities with features, we feed these additional cues through specific projection layers designed to align them with the input space of the LLM. Accordingly, the video, pose, and object features are projected into the LLM embedding space using linear projection layers $\mathcal{T}_{j}$ for each cue $j=\{v,p,o\}$, resulting in LLM input token representation of the video, pose, and object cues, respectively:
$Q_{v}=\mathcal{T}_{v}(V_{i});\hskip 14.45377ptQ_{p}=\mathcal{T}_{p}(P_{i});\hskip 14.45377ptQ_{o}=\mathcal{T}_{o}(O_{i})$ (2)
@@ 4 Experiments
4.1 Experimental Setting
Evaluation Metrics. Inspired by [20], LLVM’s ability to generate video-level descriptions is evaluated. This involves comparing the generated descriptions with ground truth and scoring them on dimensions such as Correctness of Information, Detail Orientation, Contextual Understanding, Temporal Understanding, and Consistency, with scores scaled to be bounded at $100$. Due to the subjective nature of this metric, Mementos Evaluation [51] is also conducted to assess the recognition of common action-verbs and object-nouns in the video descriptions compared to ground truth, presenting F1 scores for these classifications. However, comparing video descriptions generated by LLVMs presents a challenge due to the inherently subjective nature of these descriptions. Some objective evaluation benchmarks for LLVMs [52, 53, 54] primarily focus on video tasks involving in-the-wild activities. Therefore, this paper introduces novel benchmarks for assessing LLVM’s temporal understanding of ADL videos. We propose two new ADLMCQ benchmarks including ADLMCQ-AR and ADLMCQ-AF. ADLMCQ-AR involves multiple-choice question-answering for action recognition, where the model selects the correct action from a set of options given a question about the action performed in a video. Similarly, ADLMCQ-AF focuses on action forecasting, requiring the model to predict the next action based on the preceding actions. It is important to note that all evaluations are performed zero-shot.
@@ 4.1 Experimental Setting
Evaluation Datasets. For ADLMCQ-AR evaluation, we utilize the Charades [55] and Toyota Smarthome [56] datasets. Evaluation for ADLMCQ-AF is conducted using LEMMA [57] and Toyota Smarthome Untrimmed (TSU) [58] datasets. Video description tasks are assessed using the Charades and TSU datasets, both featuring long-duration videos with multiple actions per video. Notably, for the TSU dataset, we manually annotated video descriptions with fine-grained details regarding activities performed by elderly individuals, employing 6 human annotators for 174 videos. Our evaluation relies on these annotated descriptions, which we also provide to the community as part of the test set for ADL-X.
Table 2: Impact of ADL-X Training
Method Training ADLMCQ-AR ADLMCQ-AF Action Description (Charades)
Data (Smarthome) (LEMMA) Object Action Correctness
VideoChatGPT [20] ActivityNet 40.8 35.7 14.8 16.1 35.8
VideoChatGPT [20] NTU120 49.8 33.5 27.0 10.1 38.8
ADL-X ChatGPT [20] ADL-X 52.3 44.8 32.2 13.4 43.0
@@ 4.2 Impact of ADL-X Training on LLVMs
To understand the requirement of ADL-X, we assess VideoChatGPT [20] trained on 100K instruction pairs from ActivityNet [25], trimmed NTU120 [39], and ADL-X in Table 2. Notably, ADL-X ChatGPT, trained on ADL-X, consistently outperforms the others in both ADLMCQ-AR and ADLMCQ-AF tasks. However, it’s worth mentioning that while the baseline [20] exhibits strong performance in the action metric of Mementos, it notably underperforms in the object metric. It’s important to emphasize that ADLMCQ evaluations offer more objective and reliable assessments for understanding the temporal comprehension of LLVMs.
Table 3: Introducing Pose and Object Cues into LLAVIDAL
Method ADLMCQ-AR ADLMCQ-AF AD (Charades) AD (TSU)
Charades Smarthome LEMMA TSU Object Action Object Action
ADL-X ChatGPT 58.0 52.3 44.8 25.25 16.6 14.8 16.6 14.8
Pose QA 48.5 49.0 42.0 21.2 31.8 14.0 16.5 15.9
Pose Context (PC) 50.8 54.0 45.0 22.3 30.5 14.8 18.6 15.4
Pose Features (PF) 56.7 57.0 51.3 26.0 32.7 13.5 18.2 13.0
PC + PF 52.5 53.1 44.6 24.9 32.1 13.6 17.5 15.6
Object QA 51.1 50.1 40.3 23.0 32.1 13.7 17.0 16.0
Object Context 44.6 46.2 41.8 21.0 31.2 14.7 17.2 16.5
Object Features (OF) 59.0 58.8 52.6 27.0 33.1 14.3 18.0 17.7
PF + OF 56.2 56.1 51.0 26.6 30.4 14.1 20.0 14.1
4.3 How to introduce object and pose cues into the LLM space?
Table 3 explores the integration of pose and object cues into LLAVIDAL. We evaluate incorporating poses as QA, context (PC), and features (PF). While both pose context and features outperform the baseline ADL-X ChatGPT, projecting pose features directly into the LLM embedding space yields superior performance. This suggests the effectiveness of language contextualization for pose information. Combining pose context and features hinders performance, suggesting potential redundancy. In contrast, object cues as QA or context offer minimal discriminative information for the LLM. However, object features derived from ObjectLM significantly improve performance across most tasks, highlighting their importance in understanding ADL. A detailed analysis of these cues’ impact on ADLMCQ action classes is provided in Appendix F, revealing complementary information learned. Interestingly, LLAVIDAL with object features outperforms the model with pose features on all tasks. However, attempts to combine both pose and object features result in performance converging towards the pose-only model. We hypothesize this is due to the challenge of optimizing the projection layer $\mathcal{T}_{v}$ that effectively aligns both $\mathcal{T}_{p}$ and $\mathcal{T}_{o}$. Therefore, multi-cue integration is left for future work. Given its superior performance, LLAVIDAL with object features is used for the remainder of the paper.
@@ 4.3 How to introduce object and pose cues into the LLM space?
Table 4: Performance on Video Description. [CI: Correctness of Information, DO: Detail Orientation, CU: Contextual Understanding, TU: Temporal Understanding, Con: Consistency]
Method Training Charades TSU
Data Size Object Action CI DO CU TU Con Object Action CI DO CU TU Con
CogVLM [44] + GPT [7] 1.5B Images 19.8 9.4 44.2 42.0 33.2 33.0 40.6 16.8 6.1 41.0 37.0 37.6 34.4 40.2
CogVLM [44] + Llama [11] 1.5B Images 20.9 9.3 44.2 41.8 34.8 32.0 40.6 17.9 7.8 30.0 33.4 35.4 33.8 30.0
BLIP2 [45] + GPT [7] 1.5B Images 21.1 17.3 33.6 33.8 35.4 30.0 34.4 23.2 22.8 38.0 35.4 30.6 37.2 38.4
VideoLlama [19] 2.6M QA Pairs 14.7 15.9 32.2 32.0 36.0 34.4 39.6 21.0 13.4 33.2 30.4 31.2 34.6 42.0
VideoLlava [18] 1.2M QA Pairs 15.8 15.5 38.2 44.4 44.0 37.4 40.2 20.9 15.3 37.8 33.8 40.2 40.4 39.6
VideoChatGPT [20] 100K QA Pairs 14.8 16.1 35.8 44.2 41.6 42.2 37.8 21.8 18.0 43.0 45.8 41.4 43.0 50.0
ADL-X ChatGPT [20] 100K QA Pairs 32.2 13.4 43.0 46.8 42.2 43.8 38.6 16.6 14.8 43.0 47.2 39.6 37.6 50.0
LLAVIDAL 100K QA Pairs 33.1 14.3 51.8 54.2 44.0 49.2 41.8 18.0 17.7 46.0 48.6 42.2 45.8 58.0

### v3

@@ 5 Multi-modalities of ADL-X
HOI Features (OF) - Extracting HOI features involves two steps: (i) Action-conditioned object detection and (ii) Object Localization and Tracking. Both steps leverage off-the-shelf models that are effective without the need for additional training. Given a stitched ADL video composed of a sequence of trimmed video segments (denoted as $\textrm{clip}_{j}$), the initial step involves extracting action-conditioned object categories from each clip. This is achieved by uniformly sampling 8 frames from each video and employing a pre-trained BLIP-2 model [33] to generate a list of distinct objects observed in the frames. Subsequently, we refine this list using the ground-truth action labels and GPT-3.5. Specifically, for each $\textrm{clip}_{j}$ in a stitched video, the corresponding action label and the list of detected objects are input into GPT-3.5, which is prompted to identify the object(s) most relevant to the given action. For example, if the objects plant, chair, bottle, table are detected in a video labeled with the action Drinking, GPT-3.5 is expected to filter out and select [bottle] as the relevant object. Refer to Appendix G for our detailed action-conditioned object detection prompting strategy.
@@
$\mathbf{\mathcal{M}_{m}}$ as QA. 3D skeleton joint coordinates or relevant object trajectory coordinates are input alongside the associated action sequence into GPT-3.5 Turbo [7], which generates a general description of the skeleton motion or human-object interactions (HOI) of an ADL-X video. This description is then re-fed into GPT-3.5 Turbo to generate two QA pairs that provide detailed explanations of the skeleton and object motions. These QA pairs are then added to the training set of text queries, $Q_{t}$, to tune the LLVM instruction.
$\mathcal{M}_{m}$ as context. To integrate contextual information of human skeletons or human-object interactions, we append modality-specific information to the input text query $Q_{t}$ while training the LLVM. For $\mathcal{M}_{s}$, we initially identify five peripheral joints – the head, right hand, left hand, right knee, and left knee – due to their significant contribution to motion. For $\mathcal{M}_{o}$, we utilize the trajectory coordinates of the relevant object(s) in the videos. Using GPT-3.5 Turbo, we generate descriptions of the motion for each of these joints or objects based on their trajectories throughout the video, specifically focusing on how the joint and object coordinates evolve. The generated descriptions, denoted as $Q^{m}_{t}|m=\{s,o\}$, are subsequently appended to the text query $Q_{t}$, incorporates these skeleton or human-object descriptions as additional contextual information. This enriched query $Q_{t}^{new}=[Q^{m}_{t}\hskip 3.61371ptQ_{t}]$ is then employed for instruction tuning.
6 Multimodal Progressive (MMPro) Training
We dub our trained LLVM on on ADL-X as LLAVIDAL. Integrating QA pairs and contextual information into $Q_{t}$ is achieved using conventional training methodologies for LLVMs. However, the joint integration of time-synchronized modalities ($\mathcal{X}_{v}$, $\mathcal{X}_{s}$, and $\mathcal{X}_{o}$) presents challenges, primarily due to conflicting gradients from the different modalities. To address this, we utilize modality-specific connectors that align each modality with the LLVM input space. To mitigate the challenges of training with multiple modalities, we adopt a Multimodal Progressive (MMPro) training strategy for LLAVIDAL. This approach incrementally increases the training complexity by progressively adding modality-specific connectors, following a predefined growth schedule. These connectors project the modality-specific features into the LLVM embedding space, facilitating effective multimodal integration.
MMPro training is structured into $|\eta|$ equispaced stages, with #Total $\frac{\#\text{Total iterations}}{|\eta|}$ iterations per stage. In the case of LLAVIDAL, where we integrate three modalities via connectors into the LLM embedding space, $\eta=3$ stages. During stage 1, alignment of specific-modality with LLM embedding space is performed. Consequently, video, skeleton, and HOI features are independently projected into the LLM embedding space using linear projection layers $\mathcal{T}_{m}$ and their respective parameters $\theta_{m}$ for each cue $m=\{v,s,o\}$, resulting in LLM input token representations of the video, skeleton, and HOI cues, respectively:
$Q_{v}=\mathcal{T}_{v}(\mathcal{X}_{v};\theta_{v});\hskip 3.61371ptQ_{s}=\mathcal{T}_{s}(\mathcal{X}_{s};\theta_{s});\hskip 3.61371ptQ_{o}=\mathcal{T}_{o}(\mathcal{X}_{o};\theta_{o})\vspace{-0.1in}$ (1)
[…]
Sequence of Integrating Modalities. Our approach to determining the sequence for integrating modalities leverages the principles of curriculum learning.Curriculum learning [21], a machine learning strategy inspired by the pedagogical approach of progressing from easy to complex tasks, is utilized to structure the integration sequence in LLAVIDAL. MMPro training strategy adopts an incremental difficulty progression for multi-modal alignment. The complexity measure guiding the integration sequence is derived from the optimized autoregressive loss obtained from stage 1. Our findings suggest a gradation of integration difficulty among the modalities, with videos and skeletons being relatively simpler to align compared to HOI. The sparsity of HOI features, indicated by the infrequency of distinct HOIs within videos, explains the challenge of aligning HOI with the language embedding space. Consequently, in LLAVIDAL, the modalities are integrated in the order of skeletons followed by HOI, adhering to the curriculum of escalating difficulty.
[…]
When performing inference, LLAVIDAL utilizes only the video cue, consequently eliminating the need for person-centric cropping and additional modalities.
Figure 4: MMPro training. Our proposed three-stage progressive training pipeline used to train LLAVIDAL. Stage 1 initializes independent projections for skeleton, object, and video features. Stage 2 combines skeleton and video modalities. Stage 3 integrates all modalities into the final model. Large hollow arrows indicate weight transfer between stages.
@@ 7 Experiments
In this section we first discuss implementation details, then introduce the ADL related tasks used to evaluate state-of-the-art LLVMs, the evaluation metrics, and finally we present all of the results and analysis.
7.1 Implementation Details
During training, the embedding dimensions are $D_{v}=1024$ for visual features extracted from CLIP-L/14 [52], $D_{o}=512$ for object features extracted from OWLv2 [47], $D_{s}=216$ for skeleton features from SkeletonCLIP [63]. The embedding dimension of the Vicuna [13] LLM is $K=4096$, and the number of visual and skeleton tokens are set as $F_{v}=356$ and $F_{s}=256$ respectively. The videos input to LLAVIDAL  contain $T=100$ frames, with a spatial resolution of $H=224,W=224$. LLAVIDAL  is trained on 8 NVIDIA RTX A6000 GPUs for $3$ epochs with a total batch size of $32$ and a learning rate of $2e^{-5}$. For the purpose of promoting research in this field, we also provide the extracted skeleton features and object trajectories of LLAVIDAL along with the ADL-X dataset.
7.2 ADL Tasks, Test Datasets, and Metrics
This paper introduces novel benchmarks for assessing LLVM’s temporal understanding of diverse and real-world ADL videos. We propose two new ADL MCQ benchmarks including ADL MCQ-AR and ADL MCQ-TC. ADL MCQ-AR involves multiple-choice question-answering for action recognition, where the model selects the correct action from a set of options given a question about the action performed in a video. Similarly, ADL MCQ-TC focuses on predicting the missing action in a temporal sequence, given a video containing either a complete or incomplete sequence, including the action at the blank position (e.g., “take something off table, “drink from cup”, $<$blank$>$, “walk away”). If the $<$blank$>$ appears at the end of an incomplete sequence where the input video lacks subsequent actions, the task is an action forecasting. Finally, we also evaluate the LLVMs on ADL video description tasks for long-duration videos with densely occurring actions and fine-grained captions. Note that all our evaluations are performed zero-shot.
@@ 7.2 ADL Tasks, Test Datasets, and Metrics
Evaluation Datasets. For ADL MCQ-AR evaluation, we utilize the Charades [61] (Char) and Toyota Smarthome [18] (SH) datasets. Evaluation for ADL MCQ-TC is conducted using LEMMA [27] (LEM) and Toyota Smarthome Untrimmed [16] (TSU) datasets. ADL MCQ-TC on LEMMA is an action forecasting task. Video description tasks are assessed using the Charades and TSU datasets, both featuring long-duration videos with multiple actions per video. Notably, for the TSU dataset, we manually annotated video descriptions with fine-grained details regarding activities performed by elderly individuals, employing 6 human annotators for 174 videos. Our evaluation relies on these annotated descriptions, which we also provide to the community as part of the test set for ADL-X.
Evaluation Metrics. To ensure result consistency and reproducibility, we employ LLAMA 3.1 [66] locally instead of the GPT API, where model weights may update over time. For ADL MCQ task evaluation, we follow [42, 36, 43], prompting LLAMA 3.1 with the output of LLAVIDAL, along with MCQ questions and options, to select the nearest option generated by LLAVIDAL, thus minimizing scoring bias. For video-level description evaluation, following [45], generated descriptions are compared against ground truth and scored by LLAMA across dimensions of Information Correctness, Detail Orientation, Contextual Understanding, Temporal Understanding, and Consistency, with scores scaled to a maximum of 100.
Table 1: Impact of different components in the ADL-X data curation framework on LLVM training.
Training data PAG TS WS ADL MCQ-AR ADL MCQ-TC Descrip. (Avg)
SH TSU TSU
ActivityNet 39.6 20.9 31.9
NTU120 37.2 20.5 54.1
ADL-X 40.6 25.5 59.1
ADL-X 39.0 24.8 62.2
ADL-X 44.5 29.5 64.8
@@ 7.3 Impact of ADL-X Training on LLVMs
In Table 1, we demonstrate the effectiveness of our novel semi-automated curation of the ADL-X dataset. To validate the importance of domain-specific training, we trained an LLVM on trimmed NTU120 clips, showing that the NTU120-trained model outperforms the model trained on ActivityNet [8] for description task, while our ADL-X-trained LLVM achieves the highest performance in all ADL tasks. As shown in Table 1, all strategies—PAG, TS, and WS—used to curate ADL-X from NTU120 generate a high-quality video-instruction pairs with minimal hallucinations, enabling the LLVM to effectively understand human-centered actions and handle temporal randomness.
Verification of ADL-X. To validate the correctness of our LLM-assisted data curation pipeline, we conducted a human evaluation study where the quality of 100 video-QA pairs were rated on a scale of 1-5 based on QA pair correctness. This user study revealed an average rating of 4.1 for the the generated QA pairs, demonstrating that our three key filtering strategies successfully yields high quality video-instruction pairs.
Table 2: Strategies to integrate skeletons and HOIs in LLVM.
Method ADL MCQ-AR ADL MCQ-TC TSU
Char SH LEM TSU Descrip. (Avg)
ADL-X ChatGPT 51.0 44.5 28.6 29.5 64.8
              Video+Skeleton
Skeleton Features (SF) 52.7 42.6 33.1 30.3 66.5
Skeleton QA 47.5 43.1 25.9 29.6 61.8
Skeleton Context (SC) 48.2 45.4 27.8 30.2 66.3
SC + SF 51.2 46.2 33.5 32.5 66.8
              Video+HOI
HOI Features 53.8 48.0 32.6 37.1 68.0
HOI (YOLO+CLIP) 53.7 45.1 34.1 37.0 67.8
HOI QA 50.4 45.5 30.3 28.4 62.5
HOI Context 50.3 44.6 26.5 27.8 63.8
Table 3: Integrating all modalities in LLAVIDAL.
Method ADL MCQ-AR ADL MCQ-TC TSU
Char SH LEM TSU Descrip. (Avg)
SF + OF 53.8 40.7 32.1 33.1 65.8
X-InstructBLIP [50] 49.0 45.6 27.5 29.9 65.5
              MMPro Training
Prog. A (Token) 55.2 48.1 34.3 38.2 70.8
Prog. A (String) 52.7 45.4 32.3 34.5 65.5
Prog. B (Token) 52.8 49.4 32.8 33.0 69.2
Prog. B (String) 51.3 48.6 30.1 32.6 67.4
Prog. A (SF+SC, Token) 54.5 49.3 32.2 34.9 69.5
@@
Table 4: State-of-the-art comparison. Performance is shown on ADL MCQ and ADL Video Description tasks. [CI: Correctness of Information, DO: Detail Orientation, CU: Contextual Understanding, TU: Temporal Understanding, Con: Consistency]
Method Training ADL MCQ-AR ADL MCQ-TC Charades Avg TSU Avg
Data Size Char SH LEM TSU CI DO CU TU Con CI DO CU TU Con
CogVLM [68] + GPT [7] 1.5B Images 52.3 42.5 32.0 23.6 42 62 49.6 36.5 32.8 44.6 55.2 72.0 60.6 30.2 48.5 53.3
CogVLM [68] + Llama [66] 1.5B Images 52.8 43.2 32.5 22.5 40.2 61.8 49.5 36.5 33.5 44.3 49.8 66 56.6 29.8 40.2 48.5
BLIP2 [33] + GPT [7] 1.5B Images 50.2 39.6 28.9 20.2 39.8 60.2 47.8 36.0 37.2 44.2 48.8 66.6 63.6 45.6 39.8 52.9
VideoLlama [78] 2.6M QA Pairs 40.2 44.8 32.6 24.6 22.2 42.5 33.8 20.2 34.5 30.6 57.8 62.0 62.4 48.2 44.4 54.9
VideoLlava [38] 1.2M QA Pairs 41.8 49.2 30.0 25.5 23.6 46.4 34 20.6 33.5 31.6 30.8 54.8 42.4 30.4 44.5 40.6
VideoChatGPT [45] 100K QA Pairs 51.0 39.6 31.4 20.9 26.1 45.2 35.6 21.4 31.2 31.9 31.2 52.8 78.2 64.8 45.6 54.5
ChatUniVi [28] 3M QA Pairs 53.1 48.1 32.3 36.4 36.5 54.5 46.6 32.2 35.9 41.1 56.8 66.9 79.0 50.0 56.6 61.9
ADL-X ChatGPT [45] 100K QA Pairs 51.0 44.5 28.6 29.5 40.6 50.6 49.8 30.6 40.2 42.4 62.4 79.4 70.8 51.2 60.4 64.8
LLAVIDAL 100K QA Pairs 55.2 48.1 34.3 38.2 45.8 64.2 57.0 36.4 39.4 48.6 66.0 86.2 79.6 50.0 72.4 70.8
7.4 How to integrate modalities in LLVMs?
Table 2 explores the integration of skeleton and HOI cues alongside RGB videos in LLAVIDAL. We integrate skeleton information as features (SF), QA pairs, and context (SC) as illustrated in section 5. Both SF and SC outperform the ADL-X ChatGPT baseline (Video ChatGPT model trained with ADL-X) across most ADL tasks, with the combination of skeleton context and features (SC+SF) achieving the best results, corroborating the importance of skeleton integration for enhanced ADL understanding in LLVMs. In contrast, HOI cues used as QA or context provide limited discriminative value for the LLM. However, HOI features extracted from ObjectLM significantly improve performance across ADL tasks, showing their crucial role in ADL understanding. We find that HOI extraction using BLIP2 + OWLv2 outperforms the conventional object detector + CLIP approach, supporting our choice for extracting HOI.
Table 3 shows the performance of LLAVIDAL using various strategies to jointly integrate modalities $\mathcal{M}_{m}$, namely videos, skeletons, HOIs, and language. Joint training of all modalities and modality-specific alignment with language following X-InstructBLIP[50] underperforms the model trained in Table 2 due to difficulties in optimizing the projection layer $\mathcal{T}_{v}$ for effective alignment of both $\mathcal{T}_{p}$ and $\mathcal{T}_{o}$. In contrast, our MMPro strategy effectively retains discriminative representations from all modalities through its stage-wise integration—first video, then skeletons, and finally HOIs—which addresses optimization challenges faced by baseline multimodal models. To validate the modality integration sequence in LLAVIDAL, we conduct experiments with a curriculum-based model (Prog. A) using stage 1 MMPro losses and an alternative integration sequence—video, then objects, then skeletons (Prog. B). We also ablate the impact of prefixing each modality with a token or string. Results show that Prog. A with token-type prefixing per modality performs best across most ADL tasks, further confirming the efficacy of curriculum-based modality integration in MMPro. Moreover, we find that skeleton features alone are sufficient, and skeleton context is not needed when integrating skeletons in LLAVIDAL.